Four hundred grams of an acid solution, which is 12 percent pure acid, should be diluted
by water. How much water is needed to reduce the strength of the solution to 8 percent
pure acid?

Answer by chessace(471)

As usual, x = answer = added grams of H2O.
Currently 400*0.12 acid = 48g, 400 total.
After adding, strength = new acid (same) 48g / new total (400+x)g = 0.08
48 = 0.08 * (400+x)
x+400 = 48/0.08
x = 600-400 = 200g H2O